# FIITED: Fine-grained embedding dimension optimization during training for recommender systems

### v1

@@ Abstract
Huge embedding tables in modern Deep Learning Recommender Models (DLRM) require prohibitively large memory during training and inference. Aiming to reduce the memory footprint of training, this paper proposes FIne-grained In-Training Embedding Dimension optimization (FIITED). Given the observation that embedding vectors are not equally important, FIITED adjusts the dimension of each individual embedding vector continuously during training, assigning longer dimensions to more important embeddings while adapting to dynamic changes in data. A novel embedding storage system based on virtually-hashed physically-indexed hash tables is designed to efficiently implement the embedding dimension adjustment and effectively enable memory saving. Experiments on two industry models show that FIITED is able to reduce the size of embeddings by more than 65% while maintaining the trained model’s quality, saving significantly more memory than a state-of-the-art in-training embedding pruning method. On public click-through rate prediction datasets, FIITED is able to prune up to 93.75%-99.75% embeddings without significant accuracy loss.
1 Introduction
Huge embedding tables in modern Deep Learning Recommendation Models (DLRM) reach terabytes in size (Lian et al., 2022). Training DLRMs usually requires model parallelism (Ivchenko et al., 2022; Sethi et al., 2023), but even with embedding tables distributed over multiple compute nodes, memory still proves a scarce resource (Lian et al., 2022). Reducing the memory cost of embedding tables is crucial to enable efficient model training and deployment of DLRM and allow for sustainable model development.
The size of an embedding table is determined by the number of rows (i.e., hash size), the number of columns (i.e., embedding dimension), and the size of each value in the embedding. Cutting down any of them can lead to memory saving. This paper focuses on optimizing the embedding dimension, an open and active topic in recent recommender model research (Zheng et al., 2023). Despite recent works (Qu et al., 2022; Chen et al., 2021; Ginart et al., 2021; Liu et al., 2020; 2021; Yan et al., 2021b; Zhao et al., 2021; Zhaok et al., 2021; Xiao et al., 2022; Lyu et al., 2022; Kong et al., 2022; Yao et al., 2022), setting embedding dimensions in real-world applications still largely relies on empirical evidence, and the dimensions are usually uniform across different sparse features, which is presumably non-optimal.
Different from the normal practice, assigning suitable, non-uniform embedding dimensions is desirable for two reasons. First, typically, when a uniform dimension is used across sparse features, it is chosen to be large enough to accommodate the most information-rich features, which is an over-design to features with less information, leading to unnecessarily large memory consumption. Assigning custom embedding dimensions can naturally reduce the model size. Second, to achieve good model quality, the optimal embedding dimension is inherently non-uniform due to differences among features: exceedingly large dimensions can cause model overfitting, while too short dimensions are insufficient to express information contained in the sparse features.
@@ 1 Introduction
Based on this insight, this paper proposes FIITED, a FIne-grained In-Training Embedding Dimension optimization method, where the dimension of each embedding vector is adjusted during training. FIITED assigns more memory to more important embeddings and prunes the dimensions of less important embeddings to make better use of limited memory resources. Important embeddings are identified via importance scores computed during training which are based on embedding characteristics including access frequency and gradient norms. As a result, FIITED can reduce both the model size and the memory footprint during training to a desired amount.
@@
Compared with previous Embedding Dimension Search (EDS) methods (Qu et al., 2022; Ginart et al., 2021; Chen et al., 2021; Liu et al., 2021; Zhao et al., 2021; Xiao et al., 2022; Lyu et al., 2022; Kong et al., 2022; Yao et al., 2022), the main advantage of FIITED, as shown in Figure 1, is enabling memory saving during training by not storing embeddings at the maximum length. Reducing the memory consumption of training is highly beneficial, because it enables training huge DLRMs on hardware devices with low on-device memory (e.g., GPUs) (Wang et al., 2022), lessens the need to spend extra time transferring embeddings from off-device memory-rich hardware (e.g., CPUs and SSDs) (Sethi et al., 2022; Zhao et al., 2020), and potentially frees up memory for more features to be added to the model, which can improve the model quality. However, it is non-trivial to realize, because naively pruning embedding dimensions will result in many tiny fragments of free memory which are hard to utilize. To tackle this challenge, we propose a novel chunk-based embedding storage system.
@@
In addition to reducing the training memory footprint, FIITED brings three more advantages compared to most of the existing EDS methods. First, FIITED is fined-grained and adjusts the embedding dimension at the embedding vector level, while most EDS methods operate at the sparse feature level and sets a uniform dimension for all embeddings within the same sparse feature (Qu et al., 2022; Chen et al., 2021; Ginart et al., 2021). Fine-grained EDS is necessary because even within the same sparse feature, some embeddings may be more important than others and thus require different dimensions (Lai et al., 2023; Liu et al., 2020). Second, FIITED performs EDS during training and can take advantage of dynamic changes in data characteristics. Since the importance of an embedding vector often changes over time (Lai et al., 2023; Liu et al., 2020), it is natural to assume that their optimal dimension should vary accordingly. In-training EDS offers unique opportunities to adjust embedding dimensions over time during training, which is not allowed in pre-training or post-training EDS methods (Qu et al., 2022; Ginart et al., 2021; Chen et al., 2021; Zhao et al., 2021; Kong et al., 2022). Moreover, FIITED does not rely on any prior knowledge of the training data statistics but rather adapts to data characteristics during training, and therefore is better suited to application domains with fast changing data traits or when training data are not fully observable at the start of training, e.g., during online learning. Third, FIITED does not need any pre-training or re-training procedure, and thus has relatively low training time overhead.
@@
  • The proposed method performs fine-grained EDS, adapts to changing data characteristics over time, and does not require pre-training, re-training or prior knowledge of training data traits. Thus, FIITED is more flexible, faster and easier to use than most of the existing EDS methods, while achieving higher reduction in the model size without hurting model quality.
@@
  • Experiments on two industry models show that FIITED can reduce a significant amount of embedding size during training ($>$65% and up to 50% for the two models, respectively) while maintaining the quality of the trained model. Compared to a state-of-the-art in-training embedding pruning method, AdaEmbed (Lai et al., 2023), FIITED is able to achieve higher pruning ratios without affecting model quality. On public datasets, FIITED is able to prune up to 93.75%-99.75% embeddings on three click-through datasets and 77.5%-81.25% on one classification dataset, without significant accuracy loss.
2 Related Work
2.1 Embedding Dimension Search (EDS)
Existing EDS approaches can mainly be classified into three categories: (1) Pre-training, where embedding dimensions are decided before the actual training according to information extracted from the training dataset or a light-weight pre-training process (Qu et al., 2022; Liu et al., 2021; Ginart et al., 2021). Because the selected dimensions do not change during training, pre-training EDS misses opportunities to take advantage of changing data characteristics during training. Pre-training EDS also relies heavily on prior knowledge of the training data, which is not required by FIITED. (2) Post-training, where dimension pruning is performed after training. Additional information may be collected during training to aid pruning, and re-training is sometimes needed to boost the model performance (Chen et al., 2021; Zhao et al., 2021; Kong et al., 2022). Post-training EDS cannot reduce training memory footprint. (3) In-training, where the pruning decisions are made during training. Existing works (Liu et al., 2020; Yan et al., 2021b; Zhaok et al., 2021) add additional network structures on top of the original recommender model, e.g., new DNN layers, which help select embedding dimensions during training , but the majority of them still need to store embeddings at the maximum dimension at the time of training and thus do not reduce training memory usage. One exception is ESAPN (Liu et al., 2020), which stores embeddings at their current lengths during training by abandoning the old embeddings in memory whenever the embedding dimensions are increased. But without providing a system design, it is not clear how ESAPN can utilize the small pieces of free memory. ESAPN also does not provide control over how much memory is used during training, while FIITED can reduce the memory usage to an arbitrary desired amount. In addition, ESAPN trains one policy network per sparse feature and introduces significant training time overhead, while FIITED has negligible runtime overhead.
2.2 Embedding Compression
To decrease the size of embedding tables, there are generally four ways: (1) reduce the size of each value in the table; (2) reduce the number of rows, i.e., the hash size; (3) reduce the number of columns, i.e., the embedding dimension; and (4) deconstruct the embedding table by replacing the traditional 2D array storage format with novel designs. (1) is commonly achieved by quantization (Zhang et al., 2018; Guan et al., 2019; Yang et al., 2020). (2) involves designing novel hashing methods to reduce the number of rows while maintaining the model quality (Zhang et al., 2020; Yan et al., 2021a; Desai et al., 2022). (3) has been discussed previously in Section 2.1 and includes EDS methods performed prior to, during or after training. Unlike existing EDS methods, FIITED is able to reduce the memory consumption during training without hurting the model accuracy or incurring significant performance overhead. For (4), prior research used several novel designs for embedding storage, including constructing embeddings from two separate tables (Pansare et al., 2022), via multi-layer embeddings (Ghaemmaghami et al., 2020), and by applying transformation matrices to a small set of anchor embeddings (Liang et al., 2020).
3 Embedding Table Characteristics
In this section, we analyze characteristics of sparse features and of their embeddings, which serve as the motivation for our in-training embedding dimension optimization method. In the DLRM that we use for the analysis, each sparse feature has one or more embedding tables, and the embedding dimension is the same across all tables .
3.1 Uneven Importance of Sparse Features
Using SVD analysis as a proxy for the amount of information in each feature, we show the variation across embedding tables in Figure 2(a). For each table, we compute the number of ranks required to keep a certain percentage of squared SVD values, and plot a histogram across all tables. The histogram shows that the number of ranks varies substantially for different tables and covers a large range.
@@ 3.1 Uneven Importance of Sparse Features
Figure 2: Embedding characteristics. (a) Histogram of number of ranks needed to preserve a certain percentage of squared SVD values for embedding tables. (b) Normalized row access and gradient squared sum of randomly sampled rows from 10 biggest tables. Both axes are log-scale. (c) Importance rankings of top 10 important features change over time. (d) SVD characteristics change over time. The average absolute change is 4.4.
3.2 Uneven Importance of Embeddings
@@
Not only do vast differences exist among embedding tables, rows within the same table are not equally important either. We examine the largest 10 embedding tables in the DLRM and sample 1k rows randomly from each of them, and plot the normalized row access frequency and gradient squared sum, as shown in Figure 2(b). It can be seen that both metrics span a wide range, with a tiny fraction of rows showing significantly larger values. These “hot” rows are accessed much more frequently, resulting in a much higher gradient squared sum. Because they have been updated more, they have better quality and thus can likely contribute more to the model quality.
3.3 Change of Feature Characteristics Over Time
Besides the fact that both embedding tables and rows within the same table show significant variance, the characteristics of features and of their embeddings also change over time. In Figure 2(c), we show the difference of feature importance rank values computed over a duration of 12 days. Given a trained model, feature importance is obtained in three steps: first, remove the effect of a sparse feature $i$ by randomizing its embedding vectors; then, perform inference using the modified model; finally, quantify the importance of the sparse feature $i$ by computing the change in the model accuracy. We choose the 10 most important features on the beginning day and plot how their importance rankings change in the next 11 days. The figure shows that although the majority of the 10 features remained among the top 20, some features’ importance rankings can drop significantly to 20-40 or even above 80.
Another example of such changes can be seen in the SVD analyis of embedding tables. We compare SVD results of embedding tables trained over 2 different periods of 10 days of data that are 1 month apart. For each table, we compute the number of ranks needed to retain 80% of the squared SVD values, which is a proxy of how much information is contained in a table. We can see from Figure 2(d) that the number of ranks is quite different between the two 10-day periods, meaning that the amount of information contained in each table can change over time too, which is a result of the dynamic changes in data characteristics.
4 In-Training Embedding Dimension Optimization
In order to set appropriate embedding dimensions, one needs to capture the amount of information contained in embedding vectors, which varies across sparse features, across embeddings of the same sparse feature, and changes over time (Lai et al., 2023; Liu et al., 2020) . To tackle this challenge, we design an in-training embedding dimension optimization method that dynamically adjusts the dimension of each embedding vector during training.
4.1 FIITED Overview
Aiming to optimize embedding dimensions (i.e., dimension pruning), FIITED is built upon a state-of-the-art in-training embedding pruning method, AdaEmbed (Lai et al., 2023), which prunes entire embedding vectors during training (i.e., row pruning; each row in an embedding table stores one embedding vector). AdaEmbed assigns a row utility value to each embedding row based on row access frequency and gradient information, and the rows are pruned periodically by comparing the utilities against a pruning threshold. The threshold is decided by a pre-defined pruning ratio $p$. The current utility values are sorted in ascending order and the value at the $p$-th percentile is selected as the threshold. The value $p$ can be selected by the user and indicates the desired total size of embeddings. Rows with utilities below the threshold are pruned, while the rest remain.
To realize dimension optimization, we extend the row pruning method in AdaEmbed and divide every row in the embedding tables into K chunks, where each chunk is assigned its own pruning ratio. A utility metric is maintained during training for each chunk, and a total of K pruning thresholds are computed instead of 1, with 1 threshold per chunk. Chunks with utilities below the corresponding threshold are pruned. AdaEmbed becomes a special case of our dimension pruning method when K=1. As to how to decide the K pruning ratios, we provide two methods: (1) manual selection, and (2) dynamic generation at runtime. Details of the two methods will be explained in Section 4.2 and quantitative comparisons will be given in Section 5.
To further explain the idea, an example is illustrated in Figure 3. Each row in the diagram is an embedding vector with K embedding chunks, and each chunk has its own utility value. Given pruning ratios for the embedding chunks, a pruning threshold is computed for each chunk based on chunk utility values. For each embedding chunk, the pruning decision is made by comparing its utility against the pruning threshold. For example, in the first row, the first chunk’s utility 1.3 is bigger than the threshold 0.15, so it is kept unpruned; the last chunk has utility 1.0 which is smaller than the threshold 1.4, so it is pruned. A previously pruned chunk can be brought back if its utility becomes larger than the threshold, and the chunk’s embedding values will be re-initialized.
@@ 4.1 FIITED Overview
4.1.1 Zero padding VS. linear projection
The fine-grained dimension pruning in FIITED results in different embeddings having different lengths, which poses a challenge for the computation during training and inference. In DLRMs, dot product is usually computed among embedding vectors to capture their interaction, and it only works with embeddings of the same length. In previous works (Qu et al., 2022; Ginart et al., 2021), this problem is usually solved by inserting a Fully-Connected (FC) projection layer for each table to map embeddings of shortened lengths to the maximum length, so that any two embeddings will have the same length during dot product. But this strategy does not work for FIITED because even embeddings of the same sparse feature in the same table can have different lengths. Instead, we adopt zero padding to restore the embeddings to their original length, i.e., pruned chunks are treated as all zeros. If a fully pruned row is fetched during training, an all-zero embedding is returned.
Since computing dot products on zero-padded embeddings can cause information loss due to multiplications by zero, it may seem that padding zeros will restrict the model quality. But it is not an issue: the idea is that during training, the model can adjust itself and learn to accumulate information in the unpruned chunks. To corroborate our claim, we performed preliminary tests that compare FC layers with zero-padding during training of a Multi-Task-Multi-Label industry model. Prior to training, mixed embedding dimensions are assigned based on SVD analysis of a previously trained model, and they remain fixed during training. Model quality is measured by Normalized Entropy (NE) (He et al., 2014; Qin et al., 2020), with lower values indicating better models. Results (plotted in Figure 4) show that, compared to a baseline model with uniform embedding dimensions, zero padding only incurred a marginal NE loss (0.020%) in one of the three tasks while the other two tasks had tiny NE gains (-0.017%, -0.006%), indicating that zero-padding is indeed a feasible approach.
@@
4.2 Dimension Optimization Algorithm

### v2

@@ Abstract
Huge embedding tables in modern deep learning recommender models (DLRM) require prohibitively large memory during training and inference. This paper proposes FIITED, a system to automatically reduce the memory footprint via FIne-grained In-Training Embedding Dimension pruning. By leveraging the key insight that embedding vectors are not equally important, FIITED adaptively adjusts the dimension of each individual embedding vector during model training, assigning larger dimensions to more important embeddings while adapting to dynamic changes in data. We prioritize embedding dimensions with higher frequencies and gradients as more important. To enable efficient pruning of embeddings and their dimensions during model training, we propose an embedding storage system based on virtually-hashed physically-indexed hash tables. Experiments on two industry models and months of realistic datasets show that FIITED can reduce DLRM embedding size by more than 65% while preserving model quality, outperforming state-of-the-art in-training embedding pruning methods. On public datasets, FIITED can reduce the size of embedding tables by 2.1$\times$ to 800$\times$ with negligible accuracy drop, while improving model throughput.
I Introduction
Unlike computer vision (CV) and natural language processing (NLP) models, modern deep learning recommendation models (DLRMs) take both sparse categorical features (e.g., user IDs) and dense features (e.g., commodity price) as inputs to produce model output. In a typical deep learning recommendation model, each sparse feature is represented by an entry row in the embedding table, and each row contains the embedding vector of the sparse feature. The size of an embedding table is determined by the number of rows (i.e., hash size), the number of columns (i.e., embedding dimension size), and the size of each value in the embedding vector.
As accuracy demands increase–where improvements greater than 0.1% are deemed significant [34, 22]–modern DLRMs incorporate more sparse features and instances. This requires terabytes of memory across multiple GPUs to maintain the desired throughput [12]. Unfortunately, current model size growth outpaces hardware advancements [12], and the ever-growing embedding size can lead to the underutilization of GPU compute resources. This is because embedding operations are memory-intensive rather than computation-intensive. As such, reducing the memory cost of embedding tables is crucial for efficient DLRM execution and allows for sustainable model development.
However, determining the optimal allocation of embeddings is challenging. Sparse features, thus their associated embedding tables, vary in importance to model output, with even rows within the same table having different importance due to heterogeneous characteristics (e.g., access patterns) and dynamics over time. Moreover, using exceedingly large dimensions of embedding rows not only amplifies their memory demands but can lead to model overfitting (e.g., due to insufficient training and data), while too small dimensions are insufficient to express information contained in the sparse features. Despite recent works  [20, 1, 5, 14, 15, 26, 32, 33, 24, 16, 10, 28], setting embedding dimensions in real-world applications still heavily relies on empirical evidence, leading to great engineering effort and hard to adapt over time and across tasks.
@@ I Introduction
In this paper, we propose FIITED, a system for FIne-grained In-Training Embedding Dimension optimization. FIITED allocates more memory to more important embeddings and prunes the dimensions of less important embeddings to make better use of limited memory resources, adjusting the dimension of each embedding vector during training. Important embeddings are identified via importance scores computed during training which are based on embedding characteristics including access frequency and gradient norms. As a result, FIITED can reduce both the model size and the memory footprint during training to a desired amount.
@@
Compared with existing embedding dimension search (EDS) efforts  [20, 5, 1, 15, 32, 24, 16, 10, 28], the main advantage of FIITED, as shown in Figure 1, is enabling automated memory saving during training by not storing embeddings at the maximum length. Reducing the memory consumption of training is highly beneficial, as it enables training huge DLRMs on hardware devices with low on-device memory (e.g., GPUs) [23], lessens the need to spend extra time transferring embeddings from off-device memory-rich hardware (e.g., CPUs and SSDs) [21, 31], and potentially frees up memory for more features to be added to the model, which can improve the model quality. Moreover, by automatically identifying the right embedding dimension allocation, FIITED can reduce laboring efforts while achieving even better model accuracy. However, it is non-trivial to realize, because naively pruning embedding dimensions will result in many tiny fragments of free memory which are hard to utilize. To tackle this challenge, we propose a novel chunk-based embedding storage system.
@@
In addition to reducing the training memory footprint, FIITED brings three more advantages compared to most of the existing EDS methods. First, FIITED is fined-grained and adjusts the embedding dimension at the embedding vector level, while most EDS methods operate at the sparse feature level and set a uniform dimension for all embeddings within the same sparse feature  [20, 1, 5]. Fine-grained EDS is necessary because even within the same sparse feature, some embeddings may be more important than others and thus require different dimensions [11, 14]. Second, FIITED performs EDS during training and can take advantage of dynamic changes in data characteristics. Since the importance of an embedding vector often changes over time [11, 14], it is natural to assume that their optimal dimension should vary accordingly. In-training EDS offers unique opportunities to adjust embedding dimensions over time during training, which is not allowed in pre-training or post-training EDS methods  [20, 5, 1, 32, 10]. Moreover, FIITED does not rely on any prior knowledge of the training data statistics but rather adapts to data characteristics during training, and therefore is better suited to application domains with fast-changing data traits or when training data are not fully observable at the start of training, e.g., during online learning. Third, FIITED does not need any pre-training or re-training procedure, and thus has relatively low training time overhead.
@@
  • The proposed method performs fine-grained EDS, adapts to changing data characteristics over time, and does not require pre-training, re-training, or prior knowledge of training data traits. Thus, FIITED is more flexible, faster, and easier to use than most of the existing EDS methods, while achieving a higher reduction in the model size without hurting model quality.
@@
  • Experiments on two industry models show that FIITED can reduce a significant amount of embedding size during training ($>$65% and up to 50% for the two models, respectively) while maintaining the quality of the trained model. Compared to a state-of-the-art in-training embedding pruning method, AdaEmbed [11], FIITED is able to achieve higher pruning ratios without affecting model quality. On public datasets, FIITED is able to reduce the embedding table by 100$\times$ to 800$\times$ on three click-through-rate datasets and $\times2$ to 17.8×17.8$\times$17.8 × on one classification dataset, with little accuracy drop.
@@
Figure 2: Embedding characteristics. (a) Histogram of the number of ranks needed to preserve a certain percentage of squared SVD values for embedding tables. (b) Normalized row access and gradient squared sum of randomly sampled rows from 10 biggest tables. Both axes are log-scale. (c) Importance rankings of the top 10 important features change over time. (d) SVD characteristics change over time. The average absolute change is 4.4.
II Background and Motivations
In this section, we first briefly give an background introduction to DLRMs, and then analyze the characteristics of sparse feature embedding table observed during our experiment, which leads to the motivation of our solution.
II-A Deep Learning Recommendation Models
The input of a DLRM can be categorized into sparse features (e.g. post IDs, user IDs) and dense features (e.g. timestamps). A deep learning recommendation model (DLRM) mainly consists of two parts, an embedding table for sparse features and multiple-layer perceptron (MLP) neural networks for dense features. For dense features, the features are fed into a MLP neural network to capture the weight vectors of the features. For sparse features, the embedding vectors need to be retrieved from the embedding table first, where each row contains the embedding weight vector of a sparse feature. The retrieved embedding vectors of sparse features are then interacted (e.g. concatenated or element-wise multiplied) with the weight vectors of dense features. The interaction results are then used for downstream tasks. For example, in click-through rate prediction tasks, the result vectors are fed into another MLP neural network to produce the final prediction result. Modern DLRMs normally have a large number of sparse features as a result of the enormous number of instances. Consequently, the embedding table causes the majority of memory consumption in DLRM systems, and a reduction in the size of the embedding table can lead to a reduction in the memory footprint of the entire model.
Figure 3: Change of dimension importance on two DLRM systems with different dimension sizes. The deeper color indicates a larger importance value.
II-B Embedding Table Characteristics
In this section, we analyze the characteristics of sparse features and their embeddings, which serve as the motivation for our in-training embedding dimension optimization method. A trained private DLRM model was used for the analysis.
Heterogeneous Importance of Sparse Features. For each sparse feature’s trained embedding table, we compute the number of ranks required to keep a certain percentage of squared SVD values, which serves as a proxy of the amount of information in the table, and plot a histogram across all tables. The histogram, as seen in Figure 2(a), shows that the number of ranks varies substantially for different tables and covers a large range.
@@ II-B Embedding Table Characteristics
Heterogeneous Importance of Embeddings. Rows within the same table are not equally important either. We examine the largest 10 embedding tables in the DLRM and sample 1k rows randomly from each of them, and plot the normalized row access frequency and gradient squared sum, as shown in Figure 2(b). It can be seen that both metrics span a wide range, with a tiny fraction of rows showing significantly larger values. These “hot” rows are accessed much more frequently, resulting in a much higher gradient squared sum. Because they have been updated more, they have better quality and thus can likely contribute more to the model quality.
Heterogeneous Importance across Dimensions. Upon previous insights, we further explored the importance of embedding tables across all dimensions, where some dimensions are inherently more important than other dimensions on all entries. The importance of a specific dimension is calculated using the gradient squared sums of all embedding entries on that dimension. Figure 3 shows the change of importance of each dimension during training time, sampled from two DLRM systems with varied embedding dimensions. It can be seen that some dimensions consistently remain important, while some dimensions bear lower importance scores during the entire training time.
@@
Change of Feature Characteristics Over Time. The characteristics of features and of their embeddings also change over time. In Figure 2(c), we show how feature importance changed over a duration of 12 days. Here, the importance of a sparse feature is computed by first randomizing its embedding table and then computing how much model accuracy is decreased. We choose the 10 most important features on the beginning day and plot their importance rankings in the next 11 days. The figure shows that some features’ importance rankings changed significantly from the top 10 to 20-40 or even above 80.
Another example of such changes can be seen in the SVD analysis of embedding tables. We compare SVD results of embedding tables trained over 2 different 10-day periods of data that are 1 month apart. For each table, we compute the number of ranks needed to retain 80% of the squared SVD values, which is a proxy of how much information is contained in a table. We can see from Figure 2(d) that the number of ranks is quite different between the two 10-day periods, meaning that the amount of information contained in each table can change over time as well.
III In-Training Embedding Dimension Optimization
In order to set appropriate embedding dimensions, one needs to capture the amount of information contained in embedding vectors, which varies across sparse features, across embeddings of the same sparse feature, and changes over time [11, 14] . To tackle this challenge, we design an in-training embedding dimension optimization method that dynamically adjusts the dimension of each embedding vector during training.
III-A FIITED Overview
Aiming to optimize embedding dimensions (i.e., dimension pruning), FIITED is built upon a state-of-the-art in-training embedding pruning method, AdaEmbed [11], which prunes entire embedding vectors during training (i.e., row pruning; each row in an embedding table stores one embedding vector). AdaEmbed assigns a row utility value to each embedding row based on row access frequency and gradient information, and the rows are pruned periodically by comparing the utilities against a pruning threshold. The threshold is decided by a pre-defined pruning ratio $p$. The current utility values are sorted in ascending order and the value at the $p$-th percentile is selected as the threshold. The value $p$ can be selected by the user and indicates the desired total size of embeddings. Rows with utilities below the threshold are pruned, while the rest remain.
To realize dimension optimization, we extend the row pruning method in AdaEmbed and divide every row in the embedding tables into $K$ chunks, where each chunk is assigned its own pruning ratio. A utility metric is maintained during training for each chunk, and a total of $K$ pruning thresholds are computed instead of 1, with 1 threshold per chunk. Chunks with utilities below the corresponding threshold are pruned. AdaEmbed becomes a special case of our dimension pruning method when $K=1$. As to how to decide the $K$ pruning ratios, we provide two methods: (1) manual selection, and (2) dynamic generation at runtime. Details of the two methods will be explained in Section III-B and quantitative comparisons will be given in Section IV.
To further explain the idea, an example is illustrated in Figure 4. Each row in the diagram is an embedding vector with $K$ embedding chunks, and each chunk has its own utility value. Given pruning ratios for the embedding chunks, a pruning threshold is computed for each chunk based on chunk utility values. For each embedding chunk, the pruning decision is made by comparing its utility against the pruning threshold. For example, in the first row, the first chunk’s utility 1.3 is bigger than the threshold 0.15, so it is retained; the last chunk has utility 1.0 which is smaller than the threshold 1.4, so it is pruned. A previously pruned chunk can be brought back if its utility becomes larger than the threshold, and the chunk’s embedding values will be re-initialized.
Zero padding VS. linear projection. The fine-grained dimension pruning in FIITED results in different embeddings having different lengths, which poses a challenge for the computation during training and inference. In DLRMs, the dot product is usually computed among embedding vectors to capture their interaction, and it only works with embeddings of the same length. In previous works [20, 5], this problem is usually solved by inserting a fully connected (FC) projection layer for each table to map embeddings of shortened lengths to the maximum length, so that any two embeddings will have the same length during dot product. But this strategy does not work for FIITED because even embeddings of the same sparse feature in the same table can have different lengths. Instead, we adopt zero padding to restore the embeddings to their original length, i.e., pruned chunks are treated as all zeros. If a fully pruned row is fetched during training, an all-zero embedding is returned.
Since computing dot products on zero-padded embeddings can cause information loss due to multiplications by zero, it may seem that padding zeros will restrict the model quality. But it is not an issue: the idea is that during training, the model can adjust itself and learn to accumulate information in the unpruned chunks. To corroborate our claim, we performed preliminary tests that compare FC layers with zero-padding during training of a Multi-Task-Multi-Label industry model. Prior to training, mixed embedding dimensions are assigned based on SVD analysis of a previously trained model, and they remain fixed during training. Model quality is measured by Normalized Entropy (NE) [8, 19], with lower values indicating better models. Results (plotted in Figure 5) show that, compared to a baseline model with uniform embedding dimensions, zero padding only incurred a marginal NE loss (0.020%) in one of the three tasks while the other two tasks had tiny NE gains (-0.017%, -0.006%), indicating that zero-padding is indeed a feasible approach.
III-B Dimension Optimization Algorithm